A boat is 1000 meters from a cliff. If the angle of depression from the top of the cliff to the boat is 15 degrees, how tall is the cliff?
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

A boat is 1000 meters from a cliff. If the angle of depression from the top of the cliff to the boat is 15 degrees, how tall is the cliff?
Angle of depression = 15 degrees
The sides containing the right angle
One side is the distance and the other side is the height
Tan 15 = height / distance from cliff
tan 15 = height / 1000
Height = 1000 * tan 15
= 1000 * 0.8559
855.9 meters is the height

Answer by ikleyn(53996)

.
A boat is 1000 meters from a cliff. If the angle of depression from the top of the cliff to the boat
is 15 degrees, how tall is the cliff?
~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth is fatally wrong.
        I came to bring a correct solution.


Angle of depression = 15 degrees
The sides containing the right angle
One side is the distance and the other side is the height
tan(15°) = height / distance from cliff
tan(15°) = height / 1000
height = 1000 * tan(15°)
= 1000 * 0.26794919243
267.95 meters is the height (rounded).         ANSWER

Solved correctly.